OpenStudy (anonymous):

can you help me to simplify and write in scientific notation, (4 x 10⁵)³

OpenStudy (anonymous):

64000000000000000

OpenStudy (anonymous):

6.4e16

OpenStudy (anonymous):

6.4 × 10^16

OpenStudy (anonymous):

64

OpenStudy (anonymous):

Nope, there are 15 0's, not 15.

OpenStudy (anonymous):

I mean 16 not 15

OpenStudy (anonymous):

16

OpenStudy (unklerhaukus):

\[(a \times b)^n = a^n \times b^n\] \[(x^p)^q=x^{p\times q}\]